### v1

# Is every triangle a trajectory of an elliptical billiard?

Abstract
Using Marden’s Theorem from geometric theory of polynomials, we show that for every triangle there is a unique ellipse such that the triangle is a billiard trajectory within that ellipse. Since $3$-periodic trajectories of billiards within ellipses are examples of the Poncelet polygons, our considerations provide a new insight into the relationship between Marden’s Theorem and the Poncelet Porism, two gems of exceptional classical beauty. We also show that every parallelogram is a billiard trajectory within a unique ellipse. We prove a similar result for the self-intersecting polygonal lines consisting of two pairs of congruent sides, named “Darboux butterflies”. In each of three considered cases, we effectively calculate the foci of the boundary ellipses.
Keywords: Elliptical billiards; Marden’s Theorem; $3$-periodic billiard trajectories; $4$-periodic billiard trajectories; Darboux butterflies; conics
AMS subclass: 51N20, 30C15, 37J35, 70H06
1 Introduction
Recall that mathematical billiard in a planar domain is a dynamical system where a particle moves without constraints within the domain, and obeys the billiard reflection law when it hits the boundary [KozTrBIL, Tab2005book]. Thus, billiard trajectories are polygonal lines with vertices at the boundary, such that two consecutive sides form congruent angles with the tangent line to the boundary at their joint vertex, see Figure 1.
Figure 1: Billiard reflection law: the angle of incidence with the tangent line at the bouncing point on the boundary equals the angle of reflection.
Mathematical billiards are idealized models in many aspects: a usual billiard ball is replaced by a material point, the friction and spin are neglected. Such models have natural applications, for example in geometric optics. Here, we assume that the billiard particle is of the unit mass and it moves under the inertia between the impacts, i.e. uniformly along straight lines.
Billiards within ellipses have been intensively studied, see for example [Bolotin1990, Kozlov2003, KozTrBIL, DragRadn2011book, DragRadn2014jmd, ADSK2016, KS2018, BiMi2017, BiMi2022, Glu2021, CZ2021, GKR2021, GKR2022, DGR2022, FoVed] and references therein.
In this note, we will give the affirmative answer to the title questions. The main ingredients in our considerations of triangular trajectories are Marden’s theorem from the geometry of polynomials [MardenPOLY] and the classical Ceva Theorem from elementary geometry. Periodic trajectories of billiards within ellipses can be seen as an instance of so-called Poncelet polygons, which are closed polygonal lines inscribed in one conic an circumscribed about the other conic, see e.g. [DragRadn2011book]. The Poncelet Theorem from projective geometry of conics states that if for a given pair of conics there is one Poncelet polygon, then there are infinitely many such polygons and all have the same number of sides. Thus, we are in this work dealing with an interaction between Marden’s Theorem and the Poncelet Theorem, which are both recognized by their exceptional classical beauty. Previously, a strong relationship between these two theorems was observed in [Drag2011] in a different context.
This paper is organised as follows. We review basic facts about billiards within triangles and the Ceva Theorem in Section 2, and billiards within ellipses in Section 3.1. We recall Marden’s Theorem in Section 4. In Section 5, we prove that each triangle is a billiard trajectory within a unique ellipse. For the proof that all convex $4$-periodic elliptical billiard trajectories are parallelograms, see Section 6.1, where we also show the the converse statement, see Theorem 6.3. All nonconvex $4$-periodic elliptical billiard trajectories are described in Section 7.1 as so-called Darboux butterflies. The paper is concluded by Section 7.2, where we show that every Darboux butterfly is a billiard trajectory within a unique ellipse, see Theorem 7.12. In each of the three cases, of a triangle, a parallelogram, and a Darboux butterfly, we effectively calculate the foci of the boundary ellipse.
2 Triangles
In this section consists of two parts: in Subsection 2.1, we review main properties of the billiards within a triangle in the Euclidean plane, and in Subsection 2.2 we revisit the Ceva theorem.
2.1 Triangular billiards
A trajectory of such a billiard is a polygonal line, finite or infinite, with vertices on the sides of the triangle, such that consecutive edges of the trajectory satisfy the billiard reflection law. i.e. they form the same angle with the side of the triangle which their common vertex lie on, see Figure 2.
Figure 2: Billiard motion in a triangle.
The reflection is not well defined in the vertices of the triangle, thus we omit from our consideration trajectories falling in a vertex.
Remark 2.1
Note the following minimization property of the billiard reflection: if $X$, $Y$ are two points on the same side of the a given line $\ell$, then the length $XL+LY$, for $L\in\ell$, will be the smallest when the segments $XL$ and $LY$ satisfy the billiard reflection law off $\ell$, see Figure 3.
[…]
Figure 3: The minimization property of billiard reflection: the shortest path connecting points $X$ and $Y$ which visits line $\ell$ is the billiard path $XLY$. Its length equals the segment $XY^{\prime}$, where $Y^{\prime}$ is symmetric to $Y$ with respect to $\ell$. For any other point $L^{\prime}\in\ell$, we have $XL^{\prime}+L^{\prime}Y=XL^{\prime}+L^{\prime}Y^{\prime}>XY^{\prime}$.
One of the first natural questions for any dynamical system is to establish if its periodic trajectories exist and find them if they do. Thus, next we recall a proof of the existence a periodic trajectory within any acute triangle.
Theorem 2.2 (see e.g. Theorem 2.2 [DragRadn2011book])
Let $\triangle ABC$ be an acute triangle, and $K$, $L$, $M$ the feet of its altitudes. Then $\triangle KLM$ is the triangle with minimal perimeter inscribed in $\triangle ABC$ and it represents a unique $3$-periodic trajectory of the billiard with $\triangle ABC$.
Proof. Let $M^{\prime}$ be an arbitrary given point on side $AB$. In order to find points $K^{\prime}\in BC$, $L^{\prime}\in AC$ such that the triangle $K^{\prime}L^{\prime}M^{\prime}$ has the minimal perimeter, denote by $M_{1}$, $M_{2}$ the points symmetric to $M^{\prime}$ with respect to sides $BC$ and $AC$ respectively. Then $K^{\prime}$ and $L^{\prime}$ are intersection points of $M_{1}M_{2}$ with sides $BC$ and $AC$ respectively, see Figure 4.
[…]
Figure 4: For a fixed point $M^{\prime}$ on $AB$, we construct points $K^{\prime}$, $L^{\prime}$ on the remaining two sides of the triangle, such that $\triangle K^{\prime}L^{\prime}M^{\prime}$ has smallest possible perimeter.
The perimeter of $\triangle K^{\prime}L^{\prime}M^{\prime}$ is equal to the length of segment $M_{1}M_{2}$. We observe that $M_{1}M_{2}$ is a side of the isosceles triangle $CM_{1}M_{2}$, whose angle $\angle M_{1}CM_{2}=2\angle BCA$ does not depend on the choice of point $M^{\prime}$. Since $CM_{1}\cong CM_{2}\cong CM^{\prime}$, the segment $M_{1}M_{2}$ is shortest when the point $M^{\prime}$ is such that $CM^{\prime}$ is the altitude of the triangle $ABC$ from the vertex $A$, i.e. the minimal perimeter is attained when $M^{\prime}=M$. It immediately follows that points $K^{\prime}$ and $L^{\prime}$, as constructed above, are then also the feet of the corresponding altitudes, see Figure 5.
$A$$B$$C$$L$$K$$M$
Figure 5: The vertices of the triangle with smallest perimeter inscribed in $\triangle ABC$ are the feet $K$, $L$, $M$ of the altitudes.
The minimizing property from Remark 2.1 implies that $\triangle KLM$, as the triangle with smallest perimeter inscribed in $\triangle ABC$ is a billiard trajectory within $\triangle ABC$.
Now, let us show uniqueness. Suppose that $K_{1}L_{1}M_{1}$ is another periodic trajectory within the triangle $ABC$, such that $K_{1}\in BC$, $L_{1}\in AC$, $M_{1}\in AB$. Then, according to the billiard reflection law, we can denote: $k:=\angle CK_{1}L_{1}=\angle BK_{1}M_{1}$, $l:=\angle CL_{1}K_{1}=\angle AL_{1}M_{1}$, $m:=\angle AM_{1}L_{1}=\angle BM_{1}K_{1}$. Denoting by $\angle A$, $\angle B$, $\angle C$ the angles of the triangle $ABC$, we get that the sums of the angles in the triangles $AL_{1}M_{1}$, $BK_{1}M_{1}$, $CK_{1}L_{1}$ are:
$\angle A+l+m=\angle B+k+m=\angle C+k+l=180^{\circ},$
which gives: $k=\angle A$, $l=\angle B$, $m=\angle C$.
From there, the sides of the triangle $K_{1}L_{1}M_{1}$ are parallel to the corresponding sides of the triangle $KLM$. The assumption that $K_{1}L_{1}M_{1}$ is a periodic billiard trajectory then implies $K_{1}=K$, $L_{1}=L$, $M_{1}=M$, as shown in Figure 6. Namely, if $K_{1}$ is between $K$ and $C$, then $M_{1}$ lies in the interior of $\triangle ABC$. This follows from the similarity of triangles $KLM$ and $K_{1}L_{1}M_{1}$, by using the Thales theorem. The same argument shows that $M_{1}$ lies outside $\triangle ABC$ if $K_{1}$ is between $K$ and $B$.
$A$$B$$C$$L$$K$$M$
Figure 6: The trajectories with segments parallel to the segments of $KLM$ are $6$-periodic.
$\Box$
The previous proof essentially relies on the assumption that $\triangle ABC$ is acute, since otherwise some of the feet of the altitudes are not inner points of the sides. Thus, additional discussion in needed for the right and obtuse triangles.
It is easy to see that there are periodic billiard trajectories within a right triangle: one of them, the polygonal line $KLMNMLK$, is shown in Figure 7.
$K$$L$$M$$N$
Figure 7: A periodic trajectory of the billiard within a right triangle.
After such elementary considerations for acute and right triangles, one can stay amazed by the fact that it is not known if billiards within general obtuse triangles have any periodic trajectories! There are examples for some special cases and also an intriguing computer-based proof for the existence of periodic billiard trajectories when the obtuse angle does not exceed $100^{\circ}$, see [Sch1, Sch2].
2.2 Ceva’s theorem
Now we will formulate the classical Ceva Theorem and one of its applications.
Theorem 2.3 (Ceva’s Theorem)
Let $\triangle ABC$ be a given triangle and $O$ point not belonging to the lines $AB$, $BC$, $AC$. If the lines $AO$, $BO$, $CO$ meet the lines $BC$, $AC$, $AB$ at $K$, $L$, $M$ respectively, then
${\frac{AM}{MB}}\cdot{\frac{BK}{KC}}\cdot{\frac{CL}{LA}}=1.$ (2.1)
In Ceva’s Theorem, the signed lengths of segments are used in (2.1). For example, the quantity $AM/MB$ is positive when $M$ is between $A$ and $B$ and negative otherwise.
Notice that Ceva’s theorem can be applied to an acute triangle $\triangle ABC$ and its $3$-periodic billiard trajectory $KLM$, since the altitudes intersect at one point, see Figure 5 and Theorem 2.2.
We will use Ceva’s theorem to prove the following statement.
Lemma 2.4
Let $ABC$ be a given triangle and $K$, $L$, $M$ its feet of the altitudes from $A$, $B$, $C$ respectively. Then there exist real numbers $m_{1}$, $m_{2}$, $m_{3}$ such that point $K$ divides $BC$ in the ratio $m_{2}:m_{3}$, point $L$ divides $CA$ in the ratio $m_{3}:m_{1}$, and $M$ divides $AB$ in the ratio $m_{1}:m_{2}$. Moreover
$m_{1}:m_{2}:m_{3}=(b^{2}+c^{2}-a^{2}):(a^{2}+c^{2}-b^{2}):(a^{2}+b^{2}-c^{2}),$ (2.2)
where $a$, $b$, $c$ respectively denote the lengths of the sides $BC$, $AB$, $AC$.
Proof. Since the altitudes of a triangle intersect at the orthocenter, Ceva’s theorem implies the existence of numbers $m_{1}$, $m_{2}$, $m_{3}$. The similarities of right triangles $\triangle KBA\sim\triangle MBC$, $\triangle LCB\sim\triangle KCA$, $\triangle MAC\sim\triangle LAB$ imply the following relations: $KB/AB=MB/BC$, $CL/CB=CK/AC$, $AM/AC=LA/AB$, yielding
$m_{1}:m_{2}:m_{3}=(AL\cdot AC):(BM\cdot BA):(CK\cdot CB).$
The Pythagorean theorem gives
$CK^{2}=b^{2}-AK^{2},\quad AL^{2}=c^{2}-BL^{2},\quad BM^{2}=a^{2}-CM^{2},$
thus
$m_{1}:m_{2}:m_{3}=b\sqrt{c^{2}-BL^{2}}:c\sqrt{a^{2}-CM^{2}}:a\sqrt{b^{2}-AK^{2}}.$
Observing that $a\cdot AK=b\cdot BL=c\cdot CM=2\mathcal{A}$, where $\mathcal{A}$ is the area of the triangle, we get:
$m_{1}:m_{2}:m_{3}=\sqrt{c^{2}b^{2}-4{\mathcal{A}}^{2}}:\sqrt{a^{2}c^{2}-4{\mathcal{A}}^{2}}:\sqrt{a^{2}b^{2}-4{\mathcal{A}}^{2}}.$
Finally, applying the Heron formula $\mathcal{A}=\sqrt{s(s-a)(s-b)(s-c)}$, with $s=(a+b+c)/2$, and simplifying, we get (2.2). $\Box$
We conclude this section with the formulation of the converse Ceva’s theorem.
Theorem 2.5 (Converse Ceva’s Theorem)
If points $K$, $L$, and $M$ are chosen on the lines $BC$, $AC$ and $AB$ respectively such that the relation (2.1) is satisfied then either $AK$, $BL$ and $CM$ intersect at one point or all three are parallel.
3 Conics
3.1 Elliptical billiards
In this section, we review most important properties of billiards within an ellipse. The most famous one is the focal property of elliptic billiards, which is illustrated in Figure 8.
Proposition 3.1 (First focal property of ellipses)
Let $\mathcal{E}$ be an ellipse with foci $F_{1},F_{2}$ and $A\in\mathcal{E}$ an arbitrary point. Then segments $AF_{1}$, $AF_{2}$ satisfy the billiard reflection law off $\mathcal{E}$.
[…]
Figure 8: First focal property of ellipses.
If the billiard particle is launched from one focus of the ellipse, the focal property implies that the segments of the trajectory alternately contain the foci $F_{1}$ and $F_{2}$. Next proposition characterizes the trajectories which do not contain the foci.
Proposition 3.2
Suppose that the billiard particle is traveling along segment $s$ which does not contain any focus $F_{1}$, $F_{2}$ of the boundary ellipse $\mathcal{E}$. Then we have:
  • If both foci of $\mathcal{E}$ are on the same side of $s$, then there is unique ellipse $\mathcal{E}^{\prime}$ sharing the same foci with $\mathcal{E}$ and touching the segment. Moreover, after reflection off the boundary, the next segment of the billiard trajectory will also be tangent to $\mathcal{E}^{\prime}$, see Figure 9;
    $A$
    Figure 9: Two segments satisfying the billiard reflection law of ellipse $\mathcal{E}$ are tangent to the same conic, which is confocal with $\mathcal{E}$.
  • If $s$ crosses the segment $F_{1}F_{2}$, then there is a hyperbola $\mathcal{H}$ sharing the same foci with $\mathcal{E}$ and touching the line containing $s$. Moreover, after reflection off the boundary, the next segment of the billiard trajectory will also be tangent to $\mathcal{H}$.
Proposition 3.2 implies the following key property of elliptical billiards: each trajectory of the billiard within ellipse has a caustic – a curve such that each segment of the trajectory lies on a tangent line of that curve, see Figure 10. For elliptical billiards, caustics are ellipses hyperbolas confocal with the boundary, including degenerate ones, which can be identified as horizontal and vertical axes of the ellipse $\mathcal{E}$.
Figure 10: The caustics of billiard trajectories.
A direct consequence of Propositions 3.1 and 3.2 is the well-known second focal property of ellipses, illustrated in Figure 11.
Proposition 3.3 (Second focal property of ellipses)
Let $\mathcal{E}$ be an ellipse with foci $F_{1},F_{2}$ and $A$ an arbitrary point outside ellipse $\mathcal{E}$. Denote by $t_{1}$ and $t_{2}$ the two tangents from $A$ to ellipse $\mathcal{E}$. Then the angle between $t_{1}$ and the segment $AF_{1}$ is equal to the angle between $AF_{2}$ and $t_{2}$.
$A$
Figure 11: Second focal property of ellipses
In the next lemma, we note that the type of the caustic may be determined by the period of a trajectory.
Lemma 3.4
A periodic trajectory of the elliptic billiard with odd period has an ellipse as the caustic.
Proof. Suppose that the caustic of a given trajectory is a hyperbola. Denote by $F_{1},F_{2}$ the focal points of the boundary ellipse. Then every segment of the billiard trajectory intersects the segment $F_{1}F_{2}$. Thus, for a periodic trajectory there will be an even number of intersections of the trajectory with $F_{1}F_{2}$. Consequently, the period of a closed trajectory with hyperbola as caustic must be even. $\Box$
3.2 Pascal’s theorem
We review here the classical Pascal’s theorem.
Theorem 3.5 (Pascal Theorem, see e.g. [DragRadn2011book], Section 4.4)
Let $\mathcal{C}$ be a non-degenerate conic. Consider six points $M,N,O,P,Q,R\in\mathcal{C}$. Then the points $D=l_{MN}\cap l_{PQ}$, $E=\ell_{NO}\cap\ell_{QR}$ and $F=\ell_{OP}\cap\ell_{RM}$ are collinear.
Corollary 3.6
Let a non-degenerate conic $\mathcal{C}$ be given together with its four points $M,N,P,Q$. The intersection of the tangents to the conic at $M$ and $P$ is collinear with the points of intersection $\ell_{MN}\cap\ell_{PQ}$ and $\ell_{NP}\cap\ell_{MQ}$.
Corollary 3.7 (Simpson’s line)
Given three points $M,N,P$ on a non-degenerate conic $\mathcal{C}$. Then the intersections of the tangents to the conic at $M,N,P$ with the opposite sides of the triangle $MNP$ are collinear.
The classical Pappus theorem is a version of Pascal’s theorem for degenerate conics.
4 The Siebeck–Marden theorem
Marden’s theorem is one of the fundamental results in geometric theory of polynomials and rational functions. That theorem has a long history which is well described in the famous Marden’s book [MardenPOLY]. The earliest version of this theorem, up to our best knowledge, goes back to 1864 when Siebeck (see [Si1864]) formulated and proved it for the case of polynomials with simple roots. Here we consider only the case $n=3$.
Theorem 4.1 (Siebeck, [Si1864], $n=3$)
Let $P(z)$ be a polynomial of degree $3$ with complex coefficients, such that its zeros $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ are simple and noncollinear. Then there exists an ellipse $\mathcal{E}$ tangent to every line segment $[\alpha_{i},\alpha_{j}]$ at the midpoint, see Figure 12. Moreover, the foci of the curve $\mathcal{E}$ are zeros of the derivative polynomial $P^{\prime}(z)$.
$\beta_{1}$$\beta_{2}$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\mathcal{E}$
Figure 12: Siebeck theorem. If $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ are zeros of a cubic polynomial $P(z)$ and $\beta_{1}$, $\beta_{2}$ zeros of its derivative $P^{\prime}(z)$, then there is an ellipse with foci $\beta_{1}$, $\beta_{2}$ inscribed in triangle with vertices $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$. Moreover, the ellipse is touching the sides at midpoints.
We note that the ellipse which touches the midpoints of the triangle sides is the Steiner ellipse, i.e. the ellipse of the maximal area inscribed in that triangle. It is interesting to note that the ratio of areas of any triangle and its Steiner ellipse equals $3\sqrt{3}/{(4\pi)}$, see e.g. [GSO].
The nontrivial and interesting result of Siebeck still attracts a lot of attention, see for example [Kal2008, Drag2011]. In [Pra], this theorem is attributed to van den Berg, and two proofs are presented: the original proof of van der Berg from 1888 [Berg] and another one from [Scho]. We note that the main points of the proof of Siebeck theorem for $n=3$ from [Kal2008] are based on Propositions 3.1 and 3.2, which indicates a deep relationship with elliptic billiards. We are going to exploit that connection in next Section 5.
Siebeck’s theorem can be extended to the cases with non-simple roots, namely to the functions of the form:
$P(z)=(z-\alpha_{1})^{m_{1}}(z-\alpha_{2})^{m_{2}}(z-\alpha_{3})^{m_{3}}.$ (4.1)
Each zero of the derivative $P^{\prime}(z)$ is then either equal to some $\alpha_{i}$ such that $m_{i}>1$ or it is a zero of the logarithmic derivative of $P$:
$F(z)=(\log P(z))^{\prime}=\frac{m_{1}}{z-\alpha_{1}}+\frac{m_{2}}{z-\alpha_{2}}+\frac{m_{3}}{z-\alpha_{3}}.$ (4.2)
Notice that $F(z)$ has two zeros. The following statement then holds, see also Figure 13.
Theorem 4.2 ([MardenPOLY]*Theorem 4.2 for $n=3$)
Consider a function $P(z)$ of the form (4.1), such that $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ are noncollinear complex numbers, and $m_{1}$, $m_{2}$, $m_{3}$ non-zero real constants. Let $\beta_{1}$, $\beta_{2}$ are the zeros of the logarithmic derivative (4.2) of $P(z)$. Then there is a conic with foci $\beta_{1}$, $\beta_{2}$ touching each line of the segment $[\alpha_{i},\alpha_{j}]$ in a point dividing that segment in the ratio $m_{i}:m_{j}$.
$\beta_{1}$$\beta_{2}$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$
Figure 13: Marden theorem. An example for $m_{1}=1$, $m_{2}=2$, $m_{3}=\sqrt{3}$.
Now we are going to expand on those results from the geometric theory of polynomials, in order to give complete characterization of ellipses inscribed in triangles.
Lemma 4.3
Suppose that $F_{1}$ is a point inside a triangle. Then there exists a unique ellipse inscribed in the triangle, such that one of its foci is point $F_{1}$.
Proof. Denote the triangle by $ABC$. Denote by $F_{1}^{\prime}$, $F_{1}^{\prime\prime}$, $F_{1}^{\prime\prime\prime}$ the points symmetric to $F_{1}$ with respect to sides $BC$, $AC$, $AB$ respectively. Then the second focal point of the ellipse inscribed in $\triangle ABC$ is obtained as the intersection of bisectors of angles $F_{1}^{\prime\prime}AF_{1}^{\prime\prime\prime}$ and $F_{1}^{\prime\prime}CF_{1}^{\prime}$. Namely, if $F_{2}$ is a second focal point, then triangle $CF_{1}^{\prime\prime}F_{2}$ is congruent to triangle $CF_{1}^{\prime}F_{2}$. Thus, angle $F_{2}CF_{1}^{\prime\prime}$ is equal to angle $F_{2}CF_{1}^{\prime}$. $\Box$
Theorem 4.4
Let $ABC$ be a given triangle, and $\mathcal{E}$ an ellipse inscribed in it, touching the sides $BC$, $AC$, $AB$ in points $K$, $L$, $M$ respectively. Then the lines $AK$, $BL$, $CM$ are concurrent. (See Figure 14.)
$A$$B$$C$$M$$L$$K$
Figure 14: The lines connecting triangle vertices with the points where the inscribed ellipse touches the opposite sides are concurrent.
Proof. Let $\varphi$ be an affine transformation that maps the ellipse $\mathcal{E}$ to a circle $\mathcal{E}^{\prime}$. That circle is inscribed in triangle $\varphi(ABC)$.
Since segments of each pair $\varphi(A)\varphi(L)$ and $\varphi(A)\varphi(M)$, $\varphi(B)\varphi(K)$ and $\varphi(B)\varphi(M)$, $\varphi(C)\varphi(K)$ and $\varphi(C)\varphi(L)$ are of equal lengths as tangent segments to the circle from the same vertex, the converse Ceva’s theorem (Theorem 2.5) gives that lines $\varphi(AK)$, $\varphi(BL)$, $\varphi(CL)$ are concurrent.
The statement follows immediately. $\Box$
In order to align further this matter with the Marden theorem, we provide an alternative proof of the last theorem.
Proof. (The second proof of Theorem 4.4.) Let $F_{1}$ be one of the foci of $\mathcal{E}$. Denote the complex numbers corresponding to $A$, $B$, $C$, and $F_{1}$ by $z_{1}$, $z_{2}$, $z_{3}$, and $q$ respectively. We search for nonzero real numbers $m_{1}$, $m_{2}$, $m_{3}$ such that $q$ is one of the zeros of the derivative of $(z-z_{1})^{m_{1}}(z-z_{2})^{m_{2}}(z-z_{3})^{m_{3}}$. This imposes one complex relation on real numbers $m_{1}$, $m_{2}$, $m_{3}$. That complex relation gives two real relations which uniquely determine three real numbers $m_{1}$, $m_{2}$, $m_{3}$ up to a nonzero real factor. According to Marden’s theorem (Theorem 4.2), there is an ellipse $\mathcal{E}_{1}$ inscribed in triangle $ABC$ which has $F_{1}$ as one of its foci, with the other focus corresponding to the second zero of the derivative. Since ellipses $\mathcal{E}$ and $\mathcal{E}_{1}$ are both inscribed in $\triangle ABC$ and both have $F_{1}$ as one of their foci, Lemma 4.3 implies that they coincide. $\Box$
Theorem 4.5
Every ellipse inscribed in a triangle is a Marden ellipse, i.e. there exist positive real numbers $m_{1}$, $m_{2}$, $m_{3}$ such that the foci of the ellipse are the zeros of the logarithmic derivative of the function $(z-\alpha_{1})^{m_{1}}(z-\alpha_{2})^{m_{2}}(z-\alpha_{3})^{m_{3}}$, with $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ being the complex numbers corresponding to the vertices of the triangle. The numbers $m_{1}$, $m_{2}$, $m_{3}$ are unique up to a non-zero factor.
Proof. Denote the triangle by $ABC$, the inscribed ellipse by $\mathcal{E}$, and by $K$, $L$, $M$ the common points of $\mathcal{E}$ with sides $BC$, $AC$, $AB$ respectively. According to Theorem 4.4, the lines $AK$, $BL$, $CM$ are concurrent, thus Ceva’s theorem implies that (2.1) is satisfied. This relation determines positive numbers $m_{1}$, $m_{2}$, $m_{3}$ uniquely up to a non-zero factor. $\Box$
5 $3$-periodic trajectories of billiards within ellipses
Now, after reviewing billiards and Marden’s theory in previous sections, we are equipped and ready to address the the title question.
Theorem 5.1
Every triangle is a $3$-periodic trajectory of the billiard within an ellipse. That ellipse is uniquely determined.
Proof. Let an arbitrary triangle $KLM$ be given. Through each of its vertices, we construct a line orthogonal to the bisector of the corresponding angle of the triangle. As a result, we obtain a new triangle $ABC$, formed by these three perpendiculars. Notice that $\triangle ABC$ is acute. By construction, $KLM$ is a $3$-periodic billiard trajectory within triangle $ABC$. According to Theorem 2.2, points $K$, $L$, $M$ are the feet of the altitudes of triangle $ABC$, thus the three lines $AK$, $BL$, $CM$ intersect at one point – the orthocenter of $\triangle ABC$. Thus, according to Ceva’s Theorem (see Theorem 2.3), there exist nonzero real numbers $m_{1}$, $m_{2}$, $m_{3}$ such that point $K$ divides $BC$ in the ratio $m_{2}:m_{3}$, point $L$ divides $CA$ in the ratio $m_{3}:m_{1}$, and $M$ divides $AB$ in the ratio $m_{1}:m_{2}$. The numbers $m_{1}$, $m_{2}$, $m_{3}$ can be calculated in terms of the lengths of the sides of triangle $ABC$, see Lemma 2.4 and relation (2.2).
Now, according to Marden’s theorem (Theorem 4.2), there exists a conic $\mathcal{E}$ which touches the sides $AB$, $BC$, $CA$ respectively at $M$, $K$, $L$. Since $K$, $L$, $M$ are inner points of those sides, $\mathcal{E}$ must be an ellipse inscribed in $\triangle ABC$, see Figure 13. By construction, $KLM$ is a billiard trajectory within ellipse $\mathcal{E}$. $\Box$
6 Convex $4$-periodic trajectories of billiards within ellipses
6.1 Is every parallelogram a trajectory of an elliptical billiard?
Let us recall that every $4$-periodic convex elliptical billiard trajectory is a parallelogram, see, for example, [DR2019]*Theorem 4. In this section, in Theorem 6.3, we will show that being a convex $4$-periodic elliptical billiard trajectory represents a complete characterization of parallelograms.
We start by formulating the following useful statement.
Lemma 6.1
Let a $4$-periodic convex billiard trajectory within an ellipse be given, see Figure 15. Then we have:
  • the tangent lines to the ellipse at the reflection points form a rectangle;
  • the parallelogram is also a closed billiard trajectory within the rectangle;
  • the diagonals of the rectangle are parallel to the sides of the parallelogram.
Figure 15: A convex $4$-periodic trajectory within ellipse and the tangent lines at the points of reflection.
Proof. Follows by straightforward calculation of the angles and application of the billiard reflection law. $\Box$
Lemma 6.2
Any parallelogram is a closed billiard trajectory within a rectangle. Moreover, such a rectangle is uniquely determined for each parallelogram.
Theorem 6.3
Any parallelogram is a closed billiard trajectory within an ellipse. Moreover, such an ellipse is uniquely determined for a given parallelogram.
We immediately provide a compact, synthetic proof of this theorem, based on the Pascal theorem (see Theorem 3.5). An alternative, analytic proof is given in Section 6.2.
Proof. Let $EFGH$ be a given parallelogram. According to Lemma 6.2, there is a unique rectangle within which $EFGH$ is a closed billiard trajectory, see Figure 16.
$E$$F$$G$$H$$X$$Y$$a$$b$
Figure 16: A $4$-periodic billiard trajectory within a rectangle.
Denote by $a$ the side of the rectangle containing point $E$ and by $\mathcal{C}$ the unique conic circumscribed about $EFGH$ such that $a$ is its tangent line. We want to prove that the conic is touching the remaining sides of the rectangle exactly at the points $F$, $G$, $H$.
Applying Pascal’s theorem to $\mathcal{C}$ and its points $E$, $E$, $F$, $F$, $G$, $H$, (see Corollary 3.6) we get that the intersection points of the following pairs of lines are collinear: $a$ and $FG$; $EF$ and $GH$; $b^{\prime}$ and $AD$, where $b^{\prime}$ is the tangent line to $\mathcal{C}$ at $F$. Denote $X=a\cap FG$, $Y=b^{\prime}\cap EH$. Since $EF$ and $GH$ are parallel, this implies that the line $XY$ also must be parallel to them, thus $Y$ is the intersection point of $EH$ with the line which is parallel to $EF$ and contains $X$.
Parallelogram $EFXY$ is rhombus, since $EF=FX$, so its diagonals are orthogonal to each other, which implies that $Y\in b$, where $b$ is the side of the rectangle containing point $F$. Thus $b=b^{\prime}$.
Similarly, we can prove that the remaining two sides of the rectangle are also tangent to $\mathcal{C}$.
It remains to show that $\mathcal{C}$ is an ellipse. Each degenerate conic which contains points $E$, $F$, $G$, $H$ is intersecting transversely the sides of the rectangle, thus $\mathcal{C}$ must be non-degenerate. No two tangent lines of a parabola are parallel, thus $\mathcal{C}$ is either hyperbola or ellipse. For a hyperbola, we can notice that a pair of parallel lines cannot touch the same branch, and that the touching points of the lines containing sides of the rectangle would be placed on the extensions of those sides, which does not correspond to the geometric arrangement that we got here.
We conclude that $\mathcal{C}$ is the unique ellipse within which parallelogram $EFGH$ is a closed billiard trajectory. $\Box$
6.2 An analytic proof of Theorem 6.3
The following four lemmas and the theorem constitute an analytic proof of Theorem 6.3.
Lemma 6.4
Suppose that a rectangle and a point on one of its sides are given. Then there is a unique $4$-periodic billiard trajectory within the rectangle such that the given point is one of its vertices.
Proof. Follows from the fact that the segments of any $4$-periodic billiard trajectory within a rectangle are parallel to its diagonals, see Figure 16. $\Box$
Lemma 6.5
Consider the rectangle determined by $A,E,H$ and denote it by $AEO_{1}H$. It is similar to the rectangle $ABCD$. Let $O$ be the center of the rectangle $ABCD$. The points $A,O_{1}$ and $O$ are collinear.
Lemma 6.6
Let $F_{1},F_{2}$ be the foci of an ellipse tangent to $AB$ at $E$ and to $AD$ at $H$. Assume the foci are symmetric with respect to $O$. Denote by $F_{1}^{\prime}$ the point symmetric to $F_{1}$ w.r.t. $AB$ and by $F_{1}^{\prime\prime}$ the point symmetric to $F_{1}$ w.r.t. $AD$. Denote by $X$ and $Y$ the projections of $F_{1}$ to $AB$ and $AD$ respectively. Then:
$F_{1}^{\prime}E||OX(=\frac{1}{2}F_{1}^{\prime}F_{2}),\quad F_{1}^{\prime\prime}H||OY(=\frac{1}{2}F_{1}^{\prime\prime}F_{2})$ (6.1)
and $OX=OY$.
Note that $d=F_{1}E+F_{2}E=F_{1}^{\prime}F_{2}=F_{1}^{\prime\prime}F_{2}=2OX=2OY$, where $d$ is the defining “rope length” for the ellipse $\mathcal{E}$.
Lemma 6.7
Let $Y_{1}$ be the intersection of $EO_{1}$ and $XO$ and let $X_{1}$ be the intersection of $HO_{1}$ and $YO$. Then $AEY_{1}Y$ and $AXX_{1}H$ are rectangles.
Proof. The proof follows from the fact that $XF_{1}^{\prime}EY_{1}$ and $F_{1}^{\prime\prime}YX_{1}H$ are parallelograms. Thus, $F_{1}^{\prime}X=EY_{1}$ and $F_{1}^{\prime\prime}Y=HX_{1}$. We get $EY_{1}=XF_{1}$ and $HH_{1}=YF_{1}$. $\Box$
Theorem 6.8
Given a right angle $HAE$. Denote by $O_{1}$ the fourth vertex of the rectangle $HAEO_{1}$. Let $O$ be a point of the ray $AO_{1}$. Then, there exists a unique (up to the symmetry w.r.t. $O$) point $F_{1}$ such that its mirror images $F_{1}^{\prime}$ and $F_{1}^{\prime\prime}$ w.r.t. $AE$ and $AH$ respectively, satisfy the following
$F_{1}^{\prime}E||OX,\quad F_{1}^{\prime\prime}H||OY,\quad OX=OY,$ (6.2)
where $\{X\}=AE\cap F_{1}F_{1}^{\prime}$ and $\{Y\}=AH\cap F_{1}F_{1}^{\prime\prime}$.
Proof. Denote by $O_{x}$ and $O_{y}$ the projections of $O$ at rays $AE$ and $AH$ respectively. Denote the side lengths as follows
$AX=x,AE=e,AO_{x}=o_{x},AY=y,AH=h,AO_{y}=o_{y}.$
Here we assume the values $e,h,o_{x},o_{y}$ to be known, while $x$ and $y$ need to be determined. The condition $OX=OY$ is equivalent to
$(o_{x}-x)^{2}-(o_{y}-y)^{2}=o_{x}^{2}-o_{y}^{2},$ (6.3)
by using the Pythagorean theorem. Recall that by $X_{1}$ we denoted the intersection of $HO_{1}$ and $YO$. According to Lemma 6.7, $AXX_{1}H$ is a rectangle. Let $\{X_{2}\}=XX_{1}\cap O_{y}O$. Thus, $AXX_{2}O_{y}$ is a rectangle as well. From the similarity $\triangle OO_{y}Y\sim\triangle OX_{2}X_{1}$ we get
$(o_{x}-x)(o_{y}-y)=o_{x}(o_{y}-h).$ (6.4)
We use the notation $\mathcal{X}=o_{x}-x$ and $\mathcal{Y}=o_{y}-y$. From equations (6.3) and (6.4), we get the following biquadratic equation
${\mathcal{X}}^{4}-(o_{x}^{2}-o_{y}^{2}){\mathcal{X}}^{2}-o_{x}^{2}(o_{y}-h)^{2}=0.$ (6.5)
Let us study the corresponding quadratic equation
$Q:\hat{X}^{2}-(o_{x}^{2}-o_{y}^{2})\hat{X}-o_{x}^{2}(o_{y}-h)^{2}=0.$ (6.6)
The discriminant of (6.6) is
$\mathcal{D}_{Q}=(o_{x}^{2}-o_{y}^{2})^{2}+4o_{x}^{2}(o_{y}-h)^{2}.$
The discriminant is positive, except in the following cases:
$o_{x}=o_{y}=h,$ (6.7)
when $\mathcal{D}_{Q}=0$. Thus, quadratic equation (6.6) always has two real solutions $\hat{X}_{1},\hat{X}_{2}$. Their product
$\hat{X}_{1}\cdot\hat{X}_{2}=-o_{x}^{2}(o_{y}-h)^{2}<0,\quad o_{y}\neq h$
is negative, unless $o_{y}=h$. Thus, for $o_{y}\neq h$, exactly one of the solutions is positive, say $\hat{X}_{1}$. Thus,
$(o_{x}-x)^{2}=\hat{X}_{1}.$
From the last equation, there are two options for $x$: $x_{1}=o_{x}-\sqrt{\hat{X}_{1}}$ and $x_{2}=o_{x}+\sqrt{\hat{X}_{1}}$. One of them gives $F_{1}$ and another one $F_{2}$. Hence, $F_{1}$ and $F_{2}$ are two points symmetric with respect to $O$, the only two points that satisfy the conditions of the theorem.
The exceptional case (6.7) corresponds to the situation in which $ABCD$ is a square, with $H$ being the midpoint of $AD$. In this case the ellipse $\mathcal{E}$ is a circle and $F_{1}=F_{2}=O$.
In the case $o_{y}=h$ and $o_{x}^{2}>o_{y}^{2}$, we again get a unique positive solution $\hat{X}_{1}$.
$\Box$
Thus, we have concluded the analytic proof of Theorem 6.3. As an outcome, we get the exact formulas for the foci of the ellipse.
7 Nonconvex $4$-periodic elliptical billiard trajectories
7.1 How do nonconvex $4$-periodic trajectories look like?
Following [BLPT]*Definition 3.4, we introduce the notion of a Darboux butterfly.
Definition 7.1
Any polygonal self-intersecting line consisting of four sides, such that the first and the third are congruent as well as the second and the fourth side is called a Darboux butterfly (or a bow tie).
Lemma 7.2
Every Darboux butterfly consists of legs and diagonals of an isosceles trapezoid.
Proof. Let $GHKL$ be a Darboux butterfly. Then $\triangle GHK$ is congruent to $\triangle KLG$ and they are placed from the same side of the line $GL$. Thus, $HL$ is parallel to $GL$, and $GH$ is congruent to $KL$. $\Box$
Lemma 7.3
Every $4$-periodic elliptical billiard trajectory, having a hyperbola as a caustic, is a Darboux butterfly.
Proof. We know that every $4$-periodic convex elliptical billiard trajectory is a parallelogram, see, for example, [DR2019]*Theorem 4. From [St], Theorem 1, it follows that for each $4$-periodic trajectory of an elliptical billiard, having a hyperbola as the caustic, there exists a $4$-periodic elliptical billiard trajectory having an ellipse as the caustic such that the corresponding sides of the two trajectories are congruent. $\Box$
Thus, we come to the following question.
7.2 Is every Darboux butterfly an elliptical billiard trajectory?
We start by showing that every Darboux butterfly is a billiard trajectory within a kite.
Lemma 7.4
Every Darboux butterfly is a billiard trajectory in a unique kite, consisting of two congruent acute triangles.
Proof. Consider a Darboux butterfly $GHKL$. Denote by $S$ the intersection of $GL$ and $HK$. Let the bisector of $\angle HGS$ and the bisector of $\angle SHG$ intersect at $O_{1}$. Construct the lines $GB$ and $HB$ orthogonal to $GO_{1}$ at $G$ and to $O_{1}H$ at $H$, respectively. The angle $\angle GO_{1}H$ is obtuse since $\angle O_{1}GH+\angle O_{1}HG<\pi/2$. Since in the quadrilateral $BHO_{1}G$ the angles at $G$ and $H$ are right, we get that $\angle GBH<\pi/2$.
Let the bisector of $\angle SKL$ and the bisector of $\angle SLK$ intersect at $O_{2}$. We construct the lines $KD$ and $LD$ orthogonal to $KO_{2}$ at $K$ and to $LO_{2}$ at $L$ respectively. Denoting the intersection of $BG$ and $DK$ by $A$ and denoting the intersection of $BH$ and $DL$ by $C$, we get a quadrilateral $ABCD$, which is a kite. Polygonal line $GHKL$ is a billiard trajectory within this kite. $\Box$
Lemma 7.5
Every kite, consisting of two congruent acute triangles, has a unique Darboux butterfly $4$-periodic billiard trajectory inscribed in it. The vertices of this trajectory are the footings of the altitudes of each of these two triangles at the sides of the kite. This $4$-periodic billiard trajectory is obtained as the concatenation of two $3$-periodic billiard trajectories, each of the two being the unique $3$-periodic trajectory inscribed in each of the two acute triangles.
Proof. Consider a kite $ABCD$ with acute triangles $ABC$ and $ACD$. Let $GHKL$ be a Darboux butterfly $4$-periodic billiard trajectory in it. Let $S$ be the intersection of $GL$ and $HK$. Let $O_{1}S$ be the bisector of $\angle GSH$ and $O_{2}S$ the bisector of $\angle KSL$. Then $O_{1},S,O_{2}$ are collinear. Construct the line $\ell$ orthogonal to $O_{1}O_{2}$ at $S$. Let $\ell$ intersect $BA$ and $BC$ at $A_{1}$ and $C_{1}$, respectively. Similarly, Let $\ell$ intersect $DA$ and $DC$ at $A_{2}$ and $C_{2}$, respectively. Since $\triangle GSH\cong\triangle KLS$, the triangles $BA_{1}C_{1}$ and $DA_{2}C_{2}$ are congruent and symmetric w.r.t. $\ell$. Thus, $\ell$ coincides with $AC$ and $S\in AC$.
Since $S\in AC$ and $O_{1}O_{2}$ is orthogonal to $AC$, we see that $GSH$ and $KSL$ are $3$-periodic billiard trajectories of $\triangle ABC$ and $\triangle ADC$, respectively. Such $3$-periodic trajectories are unique within any given acute triangle and their vertices are the footings of the altitudes of the triangle, see Theorem 2.2. Thus, $G$ and $H$ are the footings of the altitudes from $A$ and $C$ in $\triangle ABC$ and $L$ and $K$ are the footings of the altitudes from $A$ and $C$ in $\triangle ADC$. $\Box$
As a small digression, we are going to show that not all $4$-periodic billiard trajectories within kites are Darboux butterflies.
Lemma 7.6
Consider a kite $ABCD$ such that the congruent triangles $\triangle ABC$ and $\triangle ADC$ are acute. Construct a line parallel to $AC$. Let $A_{1},A_{2}$, $C_{1},C_{2}$ be the intersections of that line with $AB,AD,CB$, and $CD$ respectively. Construct the footings of the altitudes of $\triangle A_{1}B_{1}C_{1}$ from $A_{1}$ and $C_{1}$, denoted as $H$ and $G$, and the altitudes from $C_{2}$ and $A_{2}$ of $\triangle DC_{2}A_{2}$, denoted $K$ and $L$. Then, $HGKL$ is a billiard trajectory.
Lemma 7.7
If an isosceles trapezoid $GKLH$ is given, then the orthogonal lines to the bisectors of its angles at the vertices form a kite $ABCD$ consisting of two congruent right triangles $ABC$ and $ADC$, with right angles at $B$ and $D$ respectively. The trapezoid is a $4$-periodic billiard trajectory within this kite.
Conversely, if a kite $ABCD$ consisting of two congruent right triangles $ABC$ and $ADC$ is given, then it admits (infinitely many) $4$-periodic billiard trajectories in the form of an isosceles trapezoid.
Theorem 7.8
Consider an acute triangle $ABC$. Let $G$ and $H$ be the footings of its altitudes from $C$ and $A$, respectively. Then, there exists a unique ellipse $\mathcal{E}$ with its foci being symmetric w.r.t. $AC$, which is tangent to $AB$ and $BC$ at $G$ and $H$, respectively.
Lemma 7.9
Consider an acute triangle $ABC$. Let $G$ and $H$ be the footings of its altitudes from $C$ and $A$ respectively. Let $\mathcal{E}$ be an ellipse with the foci being symmetric w.r.t. $AC$, which is tangent to $AB$ and $BC$ at $G$ and $H$ respectively. Let $F_{1}$ be the focal point inside $\triangle ABC$. Denote the projections of $F_{1}$ at $AB$, $BC$, and $BA$ by $O$, $Y$, and $X$ respectively. Denote by $\bar{F}_{1}"$ and $\bar{F}_{1}^{\prime}$ the projections of $F_{1}$ at the altitudes $AG$ and $CH$ respectively. Then, the points $O,\bar{F}_{1}",Y$ are collinear as well as the points $O,\bar{F}_{1}^{\prime},X$. Moreover $OX=OY$.
Proof. Let $F_{1}^{\prime}$ and $F_{1}"$ be the points symmetric to $F_{1}$ w.r.t. $AB$ and $BC$ respectively. Then, $F_{2}F_{1}^{\prime}||OX$ and $OX=F_{2}F_{1}^{\prime}/2$. The points $F_{1}^{\prime},H,F_{2}$ are collinear. The segment $\bar{F}_{1}^{\prime}H$ is parallel and congruent to $XF_{1}^{\prime}$, since $\bar{F}_{1}^{\prime}H$ and $XF_{1}^{\prime}$ are the opposite sides of a rectangle. Thus, $XF_{1}^{\prime}$ is parallel and congruent to $\bar{F}_{1}H$. We conclude that $XF_{1}H\bar{F}_{1}^{\prime}$ is a parallelogram. Thus, $XF_{1}^{\prime}$ is parallel to $F_{1}^{\prime}H$. So, $\bar{F}_{1}^{\prime}$ is collinear with $X$ and $O$. A proof that $Y,\bar{F}_{1}",O$ are collinear and $OY=F_{2}F_{1}"/2$ follows along the same lines.
Since $G$ and $H$ belong to the same ellipse with the foci $F_{1},F_{2}$, we have:
$OX=\frac{F_{2}F_{1}^{\prime}}{2}=\frac{F_{2}F_{1}"}{2}=OY.$
$\Box$
Proposition 7.10
Consider an acute triangle $ABC$. Let $G$ and $H$ be the footings of its altitudes from $C$ and $A$, respectively. There exists a unique point $F_{1}$ inside $\triangle ABC$ such that its projections $O,X,Y,\bar{F}_{1}^{\prime},\bar{F}_{1}"$ at $AC,AB,BC,CG,AH$ respectively, satisfy the following:
  • (i)
    $O,\bar{F}_{1}^{\prime},X$ are collinear;
  • (ii)
    $O,\bar{F}_{1}",Y$ are collinear;
  • (iii)
    $OX=OY$.
Lemma 7.11
Consider a triangle $ABC$ such that neither of the angles at $A$ or $C$ are right. Let $G$ be the footing of its altitudes from $C$. All the points $F$ such that the projections $O,X,\bar{F}_{1}^{\prime}$ of $F$ at $AC,AB,CG$ respectively are collinear, belong to the circle with a diameter $AC$. Conversely, any point $F$ which belongs to the circle with a diameter $AC$ has the property that the projections $O,X,\bar{F}_{1}^{\prime}$ of $F$ at $AC,AB,CG$ respectively are collinear.
Proof. We first prove Lemma 7.11. Without loss of generality, we may set $A=(0,0)$, $B=(b_{1},b_{2})$ and $C=(1,0)$. Denote the coordinates of $F=(f_{1},f_{2})$. A direct calculation shows that the collinearity condition for points $O,X,\bar{F}_{1}^{\prime}$ is equivalent to:
$CC1:\quad(b_{1}-1)b_{2}(-f_{1}+f_{1}^{2}+f_{2}^{2})=0.$ (7.1)
Since neither of the angles at $A$ or $C$ are right, $b_{1}\neq 1$ and $b_{2}\neq 0$. Thus, the condition $CC1$, (7.1), is equivalent to:
$CC1^{\prime}:\quad-f_{1}+f_{1}^{2}+f_{2}^{2}=0.$ (7.2)
The last relation (7.2) is equivalent to the condition that the point $F$ belongs to the circle
$\Big{(}f_{1}-\frac{1}{2}\Big{)}^{2}+f_{2}^{2}=\frac{1}{4},$
which is the circle with $AC$ as a diameter. $\Box$
It is interesting to note that the locus of points $F$ from Lemma 7.11 does not depend on the vertex $B$.
Proof. Now we turn to a proof of Proposition 7.10. We will use the same notation as in the proof of Lemma 7.11. From Lemma 7.11, we see that conditions (i) and (ii) are equivalent and they reduce to (7.2). A direct calculation for condition (iii) gives
$b_{2}^{2}((1-2b_{1})f_{2}^{2}+(1-b_{1})f_{1}^{2}+2(b_{1}^{2}+b_{2}^{2})f_{1}-(b_{1}^{2}+b_{2}^{2}))=0.$ (7.3)
Observing that $b_{2}\neq 0$ and substituting $f_{2}^{2}=f_{1}-f_{1}^{2}$ from (7.2) into (7.3), we get
$f_{1}=\frac{b_{1}^{2}+b_{2}^{2}}{(1-b_{1})^{2}+b_{2}^{2}}.$ (7.4)
Thus, $F_{1}=(f_{1},f_{2})$ which satisfies conditions (i), (ii), and (iii) is uniquely determined through (7.4) and (7.2). This finishes the proof of Proposition 7.10.
$\Box$
This finishes the proof of Theorem 7.8. As a consequence, we get our final
Theorem 7.12
For every Darboux butterfly there exists a unique ellipse such that the Darboux butterfly is a billiard trajectory within that ellipse.
Proof. Given a Darboux butterfly $GHKL$, according to Lemma 7.4, consider the unique kite consisting of two congruent acute triangles $ABC$ and $ADC$ for which $GHKL$ is a billiard trajectory. According to Lemma 7.5, the vertices of this trajectory are the footings of the altitudes of each of these two triangles at the sides of the kite. This $4$-periodic billiard trajectory is obtained as the concatenation of two $3$-periodic billiard trajectories, each of the two being the unique $3$-periodic trajectory inscribed in each of the two acute triangles. Thus, $G$ and $H$ are the footings of the altitudes of the acute triangle $ABC$ from $C$ and $A$ respectively. Then, according to Theorem 7.8, there exists a unique ellipse $\mathcal{E}$ with foci being symmetric w.r.t. $AC$, which is tangent to $AB$ and $BC$ at $G$ and $H$ respectively. By symmetry, this ellipse $\mathcal{E}$ is also tangent to $AD$ and $DC$ at $K$ and $L$ respectively, since $K$ and $L$ are the footings of the altitudes of the acute triangle $ADC$ from $C$ and $A$ respectively. Thus, the Darboux butterfly $GHKL$ is a billiard trajectory within the ellipse $\mathcal{E}$. The ellipse $\mathcal{E}$ is a unique conic containing four points $G,H,K,L$ and being tangent to $AB$. $\Box$
Acknowledgements
The research was supported by the Australian Research Council, Discovery Project 190101838 Billiards within quadrics and beyond, the Science Fund of Serbia grant Integrability and Extremal Problems in Mechanics, Geometry and Combinatorics, MEGIC, Grant No. 7744592 and the Simons Foundation grant no. 854861.